MODS1 Commissioning Run 6: UTC 2011 March 19

Observer(s): R. Pogge (OSU), P. Garnavich (ND), and M. Pedani (LBTO)
Telescope Operator: S. Allanson (LBTO)
AO Support: A. Rakich remote from Tucson as needed
Telescope Support: J. Urban (LBTO)
Instrument Support: D. Officer (LBTO)

Summary

Started MOS work in earnest tonight. Andrew was on for 1.5 hours to check on some active optics issues (see below). It mostly behaved itself all night except one time around UTC0843 when the primary mirror panicked (lost off ~20min). Also had a guide camera AzCamServer glitch that required us to reboot the winXP box and restart the controller starting at UTC1030 that cost about 20 minutes as well.

Winds kept us from finishing up work on the A1869 field, we had to go to targets that kept our backs to the wind. It never went over limits, however. Some clouds blew over early but otherwise very clear the rest of the night.

Tonight was of course the Supermoon (closest perigee when full), so we got an extra special does of supermoonlight to keep us on our toes. It limited some of what we could do, we only took a few spectra here and there and were mostly doing imaging. Seeing went from pretty good (0.6-arcsec) to pretty awful (3+ arcsec).

Another decent night of work.

Preparations

Paul Martini sent us an update of modsalign.py based on last-night's tests. An initial run and a couple of tweaks this afternoon shows that it is working well. All ready for tonight.

Took a new set of thru-mask lamp images at zenith for the Stone F and A1689 test masks. We'll use these to compare to the on-source thru-mask images we took the past 2 nights and (hopefully) tonight. If the mask repeatability is as good as it appears, we could well use a thru-mask image library for mask alignment, saving thru-mask images on-source for final confirmation. This could lead to better optimization, and avoids a complication of the first thru-mask images we took on-source having stars in them which complicates finding accurate centroids for the alignment holes (the stars in the holes shift the center of light, sometimes significantly, especially in dark time).

Tonight and tomorrow night the main goal is to refine our mask alignment and MOS target acquisition procedures, verifying the sign of the rotation offset, testing out the mask alignment procedures with thru-mask imaging, etc., exercising the modsalign script. Long-slit acquisition is now almost routine, and we have enough experience with it know how to make the slit acquisition tool for observers.

Full Moon tonight, so we're not going to do much spectroscopy.

Dewars filled at 1500MST, still good.

Details

All times are UTC unless otherwise noted.

0130 - Beautiful sunset tonight, but a lot of that is the colors of the mid-level cumulus clouds that started forming around us this afternoon. Most will burn off in time, but now they cover at least 30% of the sky. They are boiling up from the south, and so are likely solar powered and will be gone an hour or so after sunset. It is mostly clear overhead so it won't delay our startup. Cirrus and other clouds off to the NW where they stayed all day and all last night, so not an immediate threat.

Temperatures are still warm, 5°C at sunset, humidity 30%, winds SW at 10-15m/s. No obvious wind threat tonight, but we will be watching it.

0150 - First pointing star.
   IE = -66
   CA = 79

0153 - Collimation star - BS9122 at el=72°

A little too early of a start, high background is fuddling the GCS. Going to an acquire preset to watch it until the background drops.

0200 - now we're in business, steadily converging, seeing about 1-arcsec FWHM.

0220 - Going to BS9126 with an off-axis guide star, positions us for our first major target of the night.

Nice, seeing ~0.7-arcsec, but sometimes a little jumpy. Using a star about 3-arcmin off-axis at ~5 o'clock.

MOS Acquisiton and Mask Alignment

0230 - Going to Stone F test mask field

Seeing about 1-arcsec, we're at El=55°

Test here
  • Take an image of the field in the r filter
  • Put in the mask and take a thru-mask image
  • Compare with the mask image taken this afternoon
  • run modsalign.py with each mask image and compare

Clearly a vertical shift .

Next. Put in a +2° relative rotation, reshoot field and thru-mask image. Image mods1r.xxx.0010.fits

Result - modsalign.py is reporting rotation with the wrong sign. I modified it to report (-theta) instead of (theta).

We'll verify this after Andrew is done with the active optics work.

Active Optics Tests

0303 Doing active optics tests to see if we can replicate problems reported last night. Currently we are collimated on Stone F. Am clearing active optics to see what the dominant forces onthe mirror are when it re-collimates. Total force initial: +20

Z4 Z5 Z6 Z7 Z8 Z11 Z22
-1565.3 0 0 427 -87 0 0
-1459 0 0 150 -143 0 0
-1238 0 0 -76 -131 0 0
-639 125 -603 -217 -9 434 0
-170 -66 -317 -177 82 166 0
-60 -74 -249 -107 69 77
-38 -113 -118 -33 53 35 2
-3 -116 -79 -4 20 13 2
86 -23 -12 -9 15 -8 2
Converged but with low gain on Z22 will keep going toil the send changes sign...
66 -21 14 1 23 8 1
-5 -41 -43 30 -24 -5 3
3 6 3 9 4 16 2
7 7 1 0 -6 -1 3
33 -27 -6 -15 3 -8 3
21 18 -15 21 -20 -12 3
-59 54 8 19 -1 10 3
21 52 14 3 12 10 1
-19 -28 -11 7 -6 -8 3
52 -13 2 -8 -13 2

0330 Realize a problem with the above test; C00 removes bending forces but need to also remove the focus compensation on M1 at the same time. Also increased Z22 gain from -0.2 to -1.0.

0335 now well converged, repeating above test, total forces at outset 6.0. RX Ry at outset 110, 39.

Z4 Z5 Z6 Z7 Z8 Z11 Z22
-339 -81 -661 -22 -119 221 0
58 -109 -330 -69 34 102 0
-38 -83 -262 0 34 63 0
86 -122 -88 20 12 30 26 -6
108 -128 071 -3 43 18 0
40 8 9 34 3 25 -10
4 -23 -63 16 14 -31 5

Z2 changed sign stop here. Total corrections 3.8. Active Rx, Ry = 113, 34

0343 active preset to BS9126, collimating.
Z4 Z5 Z6 Z7 Z8 Z11 Z22
891 0 0 -276 -27 0 0
-31 -244 449 -52 -14 -187 0
-242 55 143 -24 -46 -52 0
-240 7 37 -38 -28 -11 0
-50 8 24 15 12 -34 5
-187 -40 -75 -7 -10 13 0
-115 46 -23 -11 21 -8 -3
-15 81 31 -23 27 -8 -7

converged total forces 5.6, Rx 73.9, Ry 34.6 (active optics values).

0351 sending on axis active preset to Stone D GS 0 @ 27 degrees EL

Z4 Z5 Z6 Z7 Z8 Z11 Z22
48 49 -245 168 -167 6 0
-76 26 -230 96 -67 5 0
-132 10 -151 34 -57 38 0
-258 -93 -26 33 -44 58 0
-190 -14 -186 60 -15 18 0
-130 38 -78 7 -37 10 14
-80 81 -81 12 -12 -10 14
-82 -72 -73 16 -30 10
-137 -45 -39 1 -7 34 5
-3 62 -84 30 34 -16 14
-133 -82 -15 -18 -13 48 5

collimated total corrections: -3.1, RX, 123, Ry, 72.

0400 going back to GS 0 BS9126 active on GS0. EL 78.

Z4 Z5 Z6 Z7 Z8 Z11 Z22
734 0 0 -284 306 0 0
160 -390 1125 -50 125 -159 0
43 80 403 -42 21 -90 0
-76 48 127 -55 32 -35 -14
-161 -40 33 -3 -39 4 -15
-68 -32 3 -2 24 -13 -14
-76 -34 68 4 -24 -9 -11
-24 46 -44 3 10 -6 -16
-78 -2 -2 -21 19 4 -11
-22 22 11 9 -28 -1 -10

Converged: Total correction 8.0. Rx 82. Ry. 24.

0410 presetting to Stone D GS 0, EL 23
Z4 Z5 Z6 Z7 Z8 Z11 Z22
277 36 -140 191 -153 -5 0
130 60 -197 123 -175 0 0
-39 44 -204 99 116 missed missed missed
-47 73 -163 49 -96 15 0
-35 37 -103 84 -21 30 7
1 79 -137 7 -58 36 -2
-251 -107 -118 17 -18 62 0
-170 -26 -113 38 -7 36 0
-98 12 -155 -3 -7 3 19
-223 -54 -98 -27 -11 16 0
-104 2 -69 4 -6 -3 20
-61 24 -63 8 -13 11 15

Converged: total corrections -2.2, Rx 139 Ry 86.

0420 back to GS 0 BS9126 at el 73.4

Z4 Z5 Z6 Z7 Z8 Z11 Z22
414 -321 1062 -364 429 -142 0
-16 -17 571 -121 211 -68 0
-16 -31 191 -33 49 60 68 -4
-22 42 79 -73 13 -18 -5
1 56 28 -23 40 -16 -6

converged total correction 1.0, Rx 84, Ry 10.

C00 and recollimate:

Z4 Z5 Z6 Z7 Z8 Z11 Z22
-153 -315 -393 12 -17 60 0
3 -283 -337 -26 -9 46 0
9 -121 -113 2 3 35 -9

total corrections 5. Rx 83 Ry 13

Were unable to replicate the large buildup of forces on M1, will try a similar test to this next time we see forces building up.

0433 Over to Rick.

DETXY Offset Test

0435 - Stone F field at PA=10, test of DETXY offsets Sequence:
    offsetxy 0 10 rel
    offsetxy 0 -20 rel
    offsetxy 0 10 rel
    offsetxy 0 10 abs
    offsetxy 0 -10 abs
    offsetxy 0 0 abs
Looks good. Seeiong is 1.5-arcsec and the wind is up. High clouds coming in from the west.

Stone F MOS Test Field (return to MOS acquire/align work)

0445 - Returned to the Stone F field, PA=0 to work with the mask

dialing in an artificial 1-degree rotational offset, and will see if we can recover

Pointing failed - huge offset NS (+30 arcsec?)

0455 - Pointing failed, Steven resetting
   IE = -85
   CA = +100

0458 - Resuming preset to PA=0, then offset dPA=1°

Resuming, image mods1r.xxx.0022. Cloud bands rolling over us - good news is there is nothing behind them on the satellite

0506 - Stopped by clouds - lost the guide star (actually, it first grabbed the bright bar of background at the left where the edge of the beam-splitter cube is. A shame the GCS has this orthodoxy that the hotspot has to be dead-center of the guiding subframe, but hey, what do I know, I'm just an astronomer).

Waiting for the clouds to pass and we'll start again.

Regarding the dPA=1° test interrupted by clouds, the solution from modsalign.py for the 1° introduced field rotation recovered the rotation sense, but not magnitude (1.23 instead of 1.0), and the computed dX,dY offset was larger 3-arcsec which is not believable. Will continue to poke at it, could be an issue with how the transformation is being done.

0520 - back on target

did a bunch of test images of the Stone F field, playing with the offsetting.

Target selection and slit acquisition exercise (i.e., a break from MOS transforms)

0620 - Off to another target - NGC2315 possible SNe - break from the tests on MOS and an exercise of on-the-fly target selection tools

Letting Peter fly the instrument. Seeing is 1.5-2-arcsec and bouncy. Not fun.

300s then 2x600s and a moonlight corrector. Not very pretty tonight: bright moon and bad seeing

More MOS tests - now the Abell 1689 field since Stone F is setting

0720 - BS9142 to tweak up collimation. Seeing is 1.5-2 arcsec and jumpy. Wind is up a little (we can feel the rumble)

0730 - doing the acquisition with the acqMODS and a full .acq script including slit and field image "go". Could get interesting...

Supermoon is superbright tonight, really lighting up the chamber on the webcam. It is going to be an issue here at the A1689 field.

Gathering more field and slit images and trying to align the mask, first by hand in translation, then with modsalign to get rotation.

0822 - another slit acquisition in bright moon, just briefly.

0848 - Primary mirror went down (forces were getting large). Steven is working the problem...

0914 - back to the A1689 field for mask alignment measurments and software testing. Moon is too bright to do much else tonight.

Seeing 1.6-2 arcsec, and bright moon. Doing the best we can looking at field rotation estimates using the corner stars.

0938 - wind came up suddenly, maxed out the torques on the telescope drives and we lost the guide star (we were facing into the rising wind). Aborted at image 0060. Moving out of the wind and to an azimuth where Steven can get an accurate wind speed. We'll watch that while contemplating our next moves...

0958 - Pointing to BS9162, up high in the NE where we have our backs to the wind.

1005 - converged nicely, sent to a position with a 3-arcmin off-axis star along PA=-45, rotated so star is due south of BS9162. Seeing about 1arcsec

Setting up for a PA=45 slit scan. 10sec SDSS i, 0/+20/+10/-10/-20 arcsec inside the LS60x5 fat slit.

1023 - offsetxy 0 -20 abs stalled on the TCS side. Steven is working the problem - bad polys on the rotator. Steven resetting the PCS.

1026 - resending preset. One glitch was steven forgot to load the pointing coefficients after the PCS restart. Once done we pointed fine.

1030 - during guiding, could not get image. Resent the preset. No dice. Steven restarting the GCS at 1031UTC.

1032 - preset failed. Had to restart the AzCam server on the guide camera (lbt1) after power cycling the guide camera controller from the MODS console.

1036 - Back, trying again. Nope

On the AzCamServer on lbt1, we're reading
   ERROR Readout counter stuck
   Readout aborted
as a repeated error as GCS requests an image from the guide camera.

Took these actions:
  • stopped the AzCamServer on lbt1
  • powered off the guide camera Leach controller
  • waited
  • powered on the guide camera Leach controller
  • restarted the AzCamServer on lbt1
  • restarted GCS, they shook hands...
  • resent the active preset

Same error on the first image.

now trying another more drastic step - reboot the lbt1 PC:
  • stopped the AzCamServer
  • stopped the GCS
  • restarted (rebooted) lbt1 PC
  • waiting for reboot
  • reconnect with VNC from obs2
  • AzCamServer auto restarted on reboot
  • restarted the GCS
  • resent the preset

Now working...

1048 - back to work...

Seeing was fine for a while, then blew up minutes after I started the slit alignment.

Did an offsetxy rel onto the slit, but did not "absorb", so the first sequence (images 70-75) moved back to the starting pre-offset point.

Re-did the slit centering, absorbed the offset, then ran slitscan.pro again.

That's better. Repeated to be sure (seeing got a little better after is blow up)

1123 - Rotated to PA=115 and chose another dX=0 guide star. Running the sequence again.

Seeing is better (<1 arcsec sometimes), still variable. Using z filter this time.

Got 2 good scans, mirror forces are behaving themselves now that seeing is better.

1151 - Flux Standard star, this one has a close guide star (check of probe occultation).

Target is OK, must watch for a nearby faint star 6.9arcsec NW (see direct image mods1r.20110319.0103.fits)

Spectra in 1.2 and 5-arcsec slits, 120s (could go 180 or so)

1218 - Closing up, we're done for the night.

-- RichardPogge - 18 Mar 2011
Topic revision: r11 - 16 Aug 2011, RichardPogge
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback